# Regularity Properties of Optimization-Based Controllers

### v4

@@ 1 Introduction
where $G:^{n}\times^{m}\to^{k}$ and $y\in^{k}$ denotes the output. We assume that there exists a map $h:^{m}\to^{k}$, called the steady-state map, such that for each constant input $u\in^{m}$ and initial condition $x_{0}\in^{n}$, the corresponding output of (1.2) satisfies $y(t)\to h(u)$ as $t\to\infty$. Consider the problem of driving the output to an optimal steady-state, formalized by the optimization
@@
The idea is that at each time instant, the output measurement is obtained online and fed back into (7), hence this strategy is called online feedback optimization. We note that (7) is of the form (1), by rewriting the input and output set inclusions as inequalities as in (1). In this setting, understanding the regularity properties of the closed-loop dynamics $\dot{x}=F(x,u^{*}_{\text{ofo}}(x))$ becomes necessary to ensure good performance of implementing (7) on a physical plant. Letting $u_{\text{ss}}$ denote the solution to (6), one would be interested, for instance, in showing that $u^{*}_{\text{ofo}}(x(t))\to u_{\text{ss}}$ and $y(t)\to h(u_{\text{ss}})$ as $t\to\infty$. $\bullet$
@@
Recently, projected dynamical systems have been reinterpreted from the viewpoint of control theory, to design anytime flows solving variational inequalities [36], and for understanding their relationship to controllers obtained using techniques from the theory of control barrier functions [37, 38]. For example, (11) can be interpreted as the closed-loop dynamics corresponding to the system,
@@
Even though feedback controllers of the form (13) are discontinuous, the resulting closed-loop system may still be well behaved. In the case where $\mathcal{C}$ is convex, one can show existence and uniqueness [29, Theorem 2.5] of Carathéodory solutions (cf. [39] for notions of solutions to discontinuous systems), and forward invariance of the set $\mathcal{C}$ [40, Corollary 4.8]. With the additional assumption of strong monotonicity of $H$, asymptotic stability of the unique equilibrium (11) follows as well. The control-theoretic interpretation of projected dynamical systems highlights the complex relationship between the regularity properties of optimization-based feedback controllers and the dynamical properties of the resulting closed-loop system. In particular, it shows that from the perspective of control design, a feedback controller may achieve its intended objective (e.g., ensuring invariance of a safe set or stabilization to a desired equilibrium point) with relatively weak regularity properties. $\bullet$
@@ 1.1 State of the Art
Having established the importance of characterizing the regularity properties of (1), we next discuss the state of the art. There are a variety of works in the literature [42, 43, 44, 45] and [3, Theorem 2.7] that use the theory of parametric optimization to guarantee local Lipschitz continuity or other regularity properties of optimization-based controllers. For example, the results in [42] give different conditions that ensure continuity and continuous differentiability of optimization-based controllers. However, they either require the rather strong assumption of strict complementary slackness, which is not satisfied in many cases of interest, or are limited to quadratic programs that satisfy a set of technical conditions. The paper [42] also revisits Robinson’s counterexample, first introduced in [46], in the context of optimization-based control, which shows that even for optimization problems defined by well-behaved data (e.g., second-order continuously differentiable objective function and constraints, strongly convex objective function, and feasible set with non-empty interior, which are widely employed in the design of safe and stabilizing controllers, cf. Example 1.1), the resulting controller might not be locally Lipschitz. The result in [43, Theorem 3] is more general but only ensures continuity under Slater’s condition and other regularity properties on the optimization problem. The regularity results in our previous work [44, 45] establish different Lipschitz continuity results for second-order convex programs, but are limited to this specific type of optimization problems. Finally, [3, Theorem 2.7] only guarantees continuity of optimization-based controllers derived from MPC. We also note that in some cases, $u^{*}$ can be computed in closed-form, in which case the regularity properties of $u^{*}$ can be evaluated directly without having to resort to the theory of parametric optimization. Examples of such explicitly computable optimization-based controllers are provided in [47, 12, 13, 48] and [3, Chapter 7]. We would also like to point out that even though this work is mostly focused on control laws obtained as the solution of optimization problems of the form (1), the regularity properties of other control designs has also been studied in the literature. For example, the celebrated Sontag’s Universal Formula [49] provides a smooth control law for stabilization of control-affine systems. More recently, similar designs have been given in the context of safety-critical control [50, 44, 51, 52].
@@
$f(x,\cdot)$ strongly convex $\forall x\in^{n}$
@@
$f(x,\cdot)$ strongly convex
@@
$f,g\in\mathcal{C}^{2,2}(^{n}\times^{m})$
$f(x,\cdot)$ strongly convex $\forall x\in^{n}$
@@
Nonempty feasible set
LICQ
@@
$f,g\in\mathcal{C}^{1,2}(^{n}\times^{m})$
$f(x,\cdot)$ strongly convex $\forall x\in^{n}$
@@
Nonempty feasible set
Slater’s condition
Point-Lipschitz, Hölder,
directionally differentiable
@@
$f,g\in\mathcal{C}^{0,2}(^{n}\times^{m})$
$f(x,\cdot)$ strongly convex $\forall x\in^{n}$
@@
Nonempty feasible set
Slater’s Condition
Continuous, but might
not be point-Lipschitz
@@
$f,g\in\mathcal{C}^{2,2}(^{n}\times^{m})$
$f(x,\cdot)$ strongly convex $\forall x\in^{n}$
@@ 1.2 Paper Goals and Contributions
On the technical level, the contributions of the paper are as follows. First, we show that under appropriate constraint qualifications and regularity properties of the optimization problem (1), the resulting optimization-based controller is locally Lipschitz, continuously differentiable, and even analytic. We provide specific conditions for each of these cases and observe that any of those conditions guarantee existence and uniqueness of solutions for the closed-loop system. Second, given the importance of Robinson’s counterexample in showing that optimization-based controllers defined by well-behaved data might not be locally Lipschitz and its implications, e.g., for safety-critical control, cf. Example 1.1 (where most optimization-based controllers are defined by problem data sharing the properties of Robinson’s counterexample), we characterize the regularity properties enjoyed by the parametric optimizer of problems defined by objective and constraints with the same properties as in Robinson’s counterexample. We show that even though such parametric optimizers are not locally Lipschitz in general, they enjoy other weaker regularity properties, which are enough to guarantee existence of solutions for the closed-loop system and in some special cases, even uniqueness. Third, we provide different examples that show how if the properties in Robinson’s counterexample do not hold, optimization-based controllers can be discontinuous and in some cases, even unbounded. Fourth, we show that even if the optimization-based controller is discontinuous, under appropriate regularity properties of the optimization problem (1), the parametric optimizer is measurable and locally bounded, and the closed-loop system has Filippov solutions. Finally, given a safe set of interest, we study under what regularity conditions on (1) and the set, solutions of the closed-loop system remain in the safe set, both for classical and Filippov solutions.
@@ 2.1 Notions of regularity of functions
The notion of point-Lipschitz continuity is used, for instance, in [7, Section 6.3] and called Lipschitz stability, without clearly acknowledging the difference with the notion of local Lipschitz continuity. Studying point-Lipschitz continuity is natural in the context of parametric optimization, as one is normally interested in understanding the changes in the solution with respect to a fixed value of the parameter. Locally Lipschitz functions are point-Lipschitz, but the converse is not true. For instance, the function $f:\to$ defined by $f(x)=x\sin(\frac{1}{x})$ if $x\neq 0$ and $f(0)=0$ is point-Lipschitz but not locally Lipschitz at the origin.
@@
Let $\Omega\subset^{n}$. Throughout the paper, a function $\varphi:\Omega\to^{d}$ belongs to the set $\mathcal{C}^{k}(\Omega)$ if $\varphi$ is $k$-times continuously differentiable. In case we view the elements in $\Omega$ as vectors of the Cartesian product ×${}^{m_{1}}\times^{m_{2}}$ and $\varphi$ takes the form $(x,u)\mapsto\varphi(x,u)$, the function $\varphi\in\mathcal{C}^{k,\ell}(\Omega)$ if $\varphi$ is $k$-times continuously differentiable with respect to the argument $x$ and $\ell$-times continuously differentiable with respect to the argument $u$.
@@ 2.2 Constraint Qualifications and Conditions
Linear Independence Constraint Qualification (LICQ) holds at $(x,u)\in^{n}\times^{m}$ if the vectors $\{\nabla_{u}g_{i}(x),i\in\mathcal{I}(x,u)\}$ are linearly independent;
@@ 3 Regularity of Parametric Optimizers
In this section we discuss how the assumptions on the functions $f$ and $g$ defining (1) affect the regularity properties of the resulting controller $u^{*}$. Throughout this section we assume that $f$ and $g$ belong to $\mathcal{C}^{2,2}(^{n}\times^{m})$, $f(x,\cdot)$ is strongly convex for all $x\in^{n}$, $g(x,\cdot)$ is convex for all $x\in^{n}$, and the set $\left\{u\in^{m}\;\mid\;g(x,u)\leq 0\right\}$ is nonempty for all $x\in^{n}$. Such assumptions are very common, for instance, in CBF-based QPs, (cf. Example 1.1), or in model predictive controllers for linear systems (cf. Example 1.5). Note that the convexity assumptions described above guarantee that $u^{*}(x)$ is a singleton for every $x\in^{n}$.
@@
Under the assumption that both MFCQ and CR hold at $(x,u^{*}(x))$, the parametric solution $u^{*}$ is locally Lipschitz [55, Theorem 3.6]. The same conclusion can be obtained if LICQ holds [54, Theorem 4.1] at $(x,u^{*}(x))$.
@@
Under the assumptions of LICQ and SCS, the parametric solution $u^{*}$ is continuously differentiable [53, Theorem 2.1]. This last point was already noted in the optimization-based control literature in [42, Theorem 1]. In fact, if $f$ and $g$ belong to $\mathcal{C}^{p,p}(^{n},^{m})$, with $p\in\mathbb{Z}_{>0},p\geq 2$, the proof of [53, Theorem 2.1] can be adapted using the Implicit Function Theorem for higher degree of differentiability [57, Proposition 1B.5], to show that the parametric optimizer belongs to $\mathcal{C}^{p-1}(^{n})$.
@@
Similarly, if $f$ and $g$ are analytic in n, then the proof of [53, Theorem 2.1] can be adapted using the Analytic Function Theorem [58, Theorem 3.3.2], to show that the parametric optimizer is analytic in n.
If the constraint qualifications given above do not hold, the parametric optimizer can fail to be locally Lipschitz. To illustrate this, we revisit next an example due to Robinson [46].
@@
Problem (22) is a quadratic program with strongly convex objective function, smooth objective function and constraints, and for which Slater’s condition holds for every value of the parameter (this can be shown by noting that $\hat{u}=(0,0,2+|x_{2}|,0)$ satisfies all constraints strictly). Despite these nice properties, the parametric solution of (22) is not locally Lipschitz at $(x_{1},x_{2})=(0,0)$. Indeed, let $u^{*}:^{2}\to^{4}$ denote the parametric solution of (22), with $u_{4}^{*}:^{2}\to$ its fourth component, which is given by
@@
Figure 1 depicts $u_{4}^{*}$ numerically. The other components of $u^{*}$ are continuously differentiable and therefore locally Lipschitz. However, if $p_{x_{1}}=(x_{1},\frac{1}{2}x_{1}^{2})$ and $q_{x_{1}}=(x_{1},0)$, we have
$\frac{\left\lVert u_{4}^{*}(p_{x_{1}})-u_{4}^{*}(q_{x_{1}})\right\rVert}{\left\lVert p_{x_{1}}-q_{x_{1}}\right\rVert}=\frac{1}{x_{1}}.$
Since $x_{1}$ can be taken to be arbitrarily small, this shows that $u^{*}$ is not locally Lipschitz at the origin. We also note that [55, Example 3.11] gives a similar example for a parametric quadratic program with a two-dimensional optimization variable, three-dimensional parameter, strongly convex objective function, smooth objective function and constraints, and for which Slater’s condition holds for every value of the parameter and the corresponding parametric optimizer also fails to be locally Lipschitz. $\bullet$
@@
(Regularity Properties of Parametric Optimizer): Suppose that $f$ and $g$ belong to $\mathcal{C}^{2,2}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strongly convex and $g(x,\cdot)$ is convex. Suppose that SC holds at $x_{0}\in^{n}$. Then,
@@
    there exists a neighborhood $\mathcal{V}_{x_{0}}$ of $x_{0}$ such that $u^{*}$ is point-Lipschitz at $y$ for all $y\in\tilde{\mathcal{V}}_{x_{0}}$;
@@
    $u^{*}$ has the Hölder property at $x_{0}$;
@@
    $u^{*}$ is directionally differentiable at $x_{0}$.
@@
First we note that since $f(x_{0},\cdot)$ is strongly convex and $g(x_{0},\cdot)$ is convex for all $x_{0}$, $u^{*}(x_{0})$ is unique and well-defined for all $x_{0}\in^{n}$.
To prove (i) we use [7, Theorem 6.4]. Since SC holds at $x_{0}$, by [59, Prop. 5.39], since $g(x_{0},\cdot)$ is convex, MFCQ holds at $(x_{0},u^{*}(x_{0}))$. Furthermore, since $f(x_{0},\cdot)$ is strongly convex and $g(x_{0},\cdot)$ is convex, the second-order condition SOC2 [7, Definition 6.1] holds. All of this, together with the twice continuous differentiability of $f$ and $g$ imply, by [7, Theorem 6.4], that $u^{*}$ is point-Lipschitz at $x_{0}$. Now, since $g$ is continuous, there exists a neighborhood $\mathcal{V}_{x_{0}}$ of $x_{0}$ such that SC holds for all $y\in\mathcal{V}_{x_{0}}$. By repeating the same argument, $u^{*}$ is point-Lipschitz at $y$ for all $y\in\mathcal{V}_{x_{0}}$.
Now let us prove (ii). We use [60, Theorem 2.1], which gives a sufficient condition for the solution of a variational inequality to have the Hölder property. Recall that given a map $F:^{m}\to^{m}$, and a constraint set $\bar{\mathcal{C}}\subset^{m}$, a variational inequality refers to the problem of finding $u^{*}\in\bar{\mathcal{C}}$ such that $(u-u^{*})^{T}F(u^{*})\geq 0$ for all $u\in\bar{\mathcal{C}}$. For every fixed $x\in^{n}$, by taking the map $F$ to be the gradient of $f$ with respect to $u$ at $x$, and by taking $\bar{\mathcal{C}}$ to be the constraint set of (1) at $x$, a constrained optimization problem of the form (1) can be posed as a variational inequality, cf. [61]. Since $f$ is twice continuously differentiable and strongly convex, conditions (2.1) and (2.2) in [60, Theorem 2.1] hold. Moreover, since MFCQ holds at $(x_{0},u^{*}(x_{0}))$ (because SC holds), by [62, Remark 3.6] the constraint set is pseudo-Lipschitzian [60, Definition 1.1]. All of this implies by [60, Theorem 2.1] that $u^{*}$ has the Hölder property at $x_{0}$.
Finally, (iii) follows from the fact that SC implies MFCQ and [63, Theorem 1]. ∎
@@
Figure 1: Numerical depiction of the fourth component of the parametric optimizer of Robinson’s counterexample, cf. (22). The plot shows that it is continuous at the origin, in agreement with Proposition 3.2.
@@
(Scalar parametric quadratic programs have locally Lipschitz optimizers): Suppose that $f\in\mathcal{C}^{2,2}(^{n}\times^{m})$ and and for $i\in\{1,\ldots,p\}$, let $g_{i}^{0}:^{n}\to$, $g_{i}^{1}:^{n}\to$ belong to $\mathcal{C}^{2}(^{n})$ and

### v5

@@ 1 Introduction
where $G:^{n}\times^{m}\to^{k}$ and $y\in^{k}$ denote the output. We assume that there exists a map $h:^{m}\to^{k}$, called the steady-state map, such that for each constant input $u\in^{m}$ and initial condition $x_{0}\in^{n}$, the corresponding output of (1.2) satisfies $y(t)\to h(u)$ as $t\to\infty$. Consider the problem of driving the output to an optimal steady-state, formalized by the optimization
@@
The idea is that at each time instant, the output measurement is obtained online and fed back into (7), hence this strategy is called online feedback optimization. We note that (7) is of the form (1), by rewriting the input and output set inclusions as inequalities as in (1). In this setting, understanding the regularity properties of the closed-loop dynamics $\dot{x}=F(x,u^{*}_{\text{ofo}}(x))$ becomes necessary to ensure good performance of implementing (7) on a physical plant. Letting $u_{\text{ss}}$ denote the solution to (6), one would be interested, for instance, in showing that $u^{*}_{\text{ofo}}(x(t))\to u_{\text{ss}}$ and $y(t)\to h(u_{\text{ss}})$ as $t\to\infty$. We also note that if the cost function in (5) is time-varying, the resulting “dynamic” formulation (7) is also time-varying. However, the added time-dependence can be treated as an extra parameter in the optimization problem (7), and therefore the results outlined in this paper can also be applied in such time-dependent settings. $\bullet$
@@
Recently, projected dynamical systems have been reinterpreted from the viewpoint of control theory, to design anytime flows solving variational inequalities [36], and for understanding their relationship to controllers obtained using techniques from the theory of control barrier functions [37]. For example, (11) can be interpreted as the closed-loop dynamics corresponding to the system,
@@
Even though feedback controllers of the form (13) are discontinuous, the resulting closed-loop system may still be well behaved. In the case where $\mathcal{C}$ is convex, one can show existence and uniqueness [29, Theorem 2.5] of Carathéodory solutions (cf. [38] for notions of solutions to discontinuous systems), and forward invariance of the set $\mathcal{C}$ [39, Corollary 4.8]. With the additional assumption of strong monotonicity of $H$, asymptotic stability of the unique equilibrium (11) follows as well. The control-theoretic interpretation of projected dynamical systems highlights the complex relationship between the regularity properties of optimization-based feedback controllers and the dynamical properties of the resulting closed-loop system. In particular, it shows that from the perspective of control design, a feedback controller may achieve its intended objective (e.g., ensuring invariance of a safe set or stabilization to a desired equilibrium point) with relatively weak regularity properties. $\bullet$
@@ 1.1 State of the Art
Having established the importance of characterizing the regularity properties of (1), we next discuss the state of the art. There are a variety of works in the literature [41, 42, 43, 44] and [3, Theorem 2.7] that use the theory of parametric optimization to guarantee local Lipschitz continuity or other regularity properties of optimization-based controllers. For example, the results in [41] give different conditions that ensure continuity and continuous differentiability of optimization-based controllers. However, they either require the rather strong assumption of strict complementary slackness, which is not satisfied in many cases of interest, or are limited to quadratic programs that satisfy a set of technical conditions. The paper [41] also revisits Robinson’s counterexample, first introduced in [45], in the context of optimization-based control, which shows that even for optimization problems defined by well-behaved data (e.g., second-order continuously differentiable objective function and constraints, strongly convex objective function, and feasible set with non-empty interior, which are widely employed in the design of safe and stabilizing controllers, cf. Example 1.1), the resulting controller might not be locally Lipschitz. The result in [42, Theorem 3] is more general but only ensures continuity under Slater’s condition and other regularity properties on the optimization problem. The regularity results in our previous work [43, 44] establish different Lipschitz continuity results for second-order convex programs, but are limited to this specific type of optimization problems. Finally, [3, Theorem 2.7] only guarantees continuity of optimization-based controllers derived from MPC. We also note that in some cases, $u^{*}$ can be computed in closed-form, in which case the regularity properties of $u^{*}$ can be evaluated directly without having to resort to the theory of parametric optimization. Examples of such explicitly computable optimization-based controllers are provided in [46, 12, 13, 47] and [3, Chapter 7]. We would also like to point out that even though this work is mostly focused on control laws obtained as the solution of optimization problems of the form (1), the regularity properties of other control designs has also been studied in the literature. For example, the celebrated Sontag’s Universal Formula [48] provides a smooth control law for stabilization of control-affine systems. More recently, similar designs have been given in the context of safety-critical control [49, 43, 50, 51].
@@
$f(x,\cdot)$ strictly convex $\forall x\in^{n}$
@@
$f(x,\cdot)$ strictly convex
@@
$f,g\in\mathcal{C}^{2}(^{n}\times^{m})$
$f(x,\cdot)$ strictly convex $\forall x\in^{n}$
@@
Existence of minimizer, LICQ
@@
$f,g\in\mathcal{C}^{2}(^{n}\times^{m})$
$f(x,\cdot)$ strictly convex $\forall x\in^{n}$
@@
Existence of minimizer, SC
Point-Lipschitz and Hölder,
@@
$f,g\in\mathcal{C}^{2}(^{n}\times^{m})$
$f(x,\cdot)$ strictly convex $\forall x\in^{n}$
@@
Existence of minimizer, SC
Directionally differentiable,
cf. Proposition 3.2,
locally Lipschitz for scalar
QPs, cf. Proposition 3.3, and
continuous, cf. [55, Thm 5.3],
but might not be
point-Lipschitz,
@@
cf. Example 4.2
$f,g\in\mathcal{C}^{0}(^{n}\times^{m})$
$f(x,\cdot)$ strictly convex $\forall x\in^{n}$
@@ 1.2 Paper Goals and Contributions
On the technical level, the contributions of the paper are as follows. First, we show that under appropriate constraint qualifications and regularity properties of the optimization problem (1), the resulting optimization-based controller is continuous, locally Lipschitz, continuously differentiable, and even analytic. We provide specific conditions for each of these cases and observe that any of those conditions guarantee existence and uniqueness of solutions for the closed-loop system. Second, given the importance of Robinson’s counterexample in showing that optimization-based controllers defined by well-behaved data might not be locally Lipschitz and its implications, e.g., for safety-critical control, cf. Example 1.1 (where most optimization-based controllers are defined by problem data sharing the properties of Robinson’s counterexample), we characterize the regularity properties enjoyed by the parametric optimizer of problems defined by objective and constraints with the same properties as in Robinson’s counterexample. We show that even though such parametric optimizers are not locally Lipschitz in general, they enjoy other weaker regularity properties, which are enough to guarantee existence of solutions for the closed-loop system and in some special cases, even uniqueness. Third, we provide different examples that show how if the properties in Robinson’s counterexample do not hold, optimization-based controllers can be discontinuous and in some cases, even unbounded. Fourth, we show that even if the optimization-based controller is discontinuous, under appropriate regularity properties of the optimization problem (1), the parametric optimizer is measurable and locally bounded, and the closed-loop system has Filippov solutions. Finally, given a safe set of interest, we study under what regularity conditions on (1) and the set, solutions of the closed-loop system remain in the safe set, both for classical and Filippov solutions.
@@ 2.1 Notions of regularity of functions
The notion of point-Lipschitz continuity is used, for instance, in [7, Section 6.3] and called Lipschitz stability, without clearly acknowledging the difference with the notion of local Lipschitz continuity. In the optimization literature, this property is sometimes referred to as calmness (cf. [56, Chapter 8.F]). Studying point-Lipschitz continuity is natural in the context of parametric optimization, as one is normally interested in understanding the changes in the solution with respect to a fixed value of the parameter. Locally Lipschitz functions are point-Lipschitz, but the converse is not true. For instance, the function $f:\to$ defined by $f(x)=x\sin(\frac{1}{x})$ if $x\neq 0$ and $f(0)=0$ is point-Lipschitz but not locally Lipschitz at the origin.
@@
Let $\Omega\subset^{n}$. Throughout the paper, a function $\varphi:\Omega\to^{d}$ belongs to the set $\mathcal{C}^{k}(\Omega)$ if $\varphi$ is $k$-times continuously differentiable in $\Omega$. A function $\varphi:\Omega\to^{d}$ belongs to the set $\mathcal{C}^{0}(\Omega)$ if $\varphi$ is continuous in $\Omega$. In case we view the elements in $\Omega$ as vectors of the Cartesian product ×${}^{m_{1}}\times^{m_{2}}$ and $\varphi$ takes the form $(x,u)\mapsto\varphi(x,u)$, the function $\varphi\in\mathcal{C}^{0,k}(\Omega)$ if for every $x\in\Omega$, the derivatives of order up to $k$ of $\varphi(x,\cdot)$ with respect to $u$ exist and are continuous with respect to $x$ and $u$.
@@ 2.2 Constraint Qualifications and Conditions
Linear Independence Constraint Qualification (LICQ) holds at $(x,u)\in^{n}\times^{m}$ if the vectors $\{\nabla_{u}g_{i}(x,u),i\in\mathcal{I}(x,u)\}$ are linearly independent;
@@ 3 Regularity of Parametric Optimizers
In this section we discuss how the assumptions on the functions $f$ and $g$ defining (1) affect the regularity properties of the resulting controller $u^{*}$. Throughout this section, we assume that $f$ and $g$ belong to $\mathcal{C}^{2}(^{n}\times^{m})$, $f(x,\cdot)$ is strictly convex for all $x\in^{n}$ (for some specific results, we further assume that $f(x,\cdot)$ is strongly convex for all $x\in^{n}$, but we make it explicit if this is the case) and $g(x,\cdot)$ is convex for all $x\in^{n}$. We further assume that, for each $x\in^{n}$, (1) has at least one minimizer (note that if $f(x,\cdot)$ is strongly convex for all $x\in^{n}$, this holds if (1) is feasible for all $x\in^{n}$). By the convexity assumptions described above, this implies that $u^{*}(x)$ is a singleton for every $x\in^{n}$. Furthermore, the assumptions on convexity also ensure that for each $x\in^{n}$, interior-point algorithms [58] can be used to solve (1), which run in polynomial time when the optimization problem is a linear program, a quadratic program, a second-order convex program, or a semidefinite program. This type of assumptions are very common, for instance, in CBF-based QPs, (cf. Example 1.1), or in model predictive controllers for linear systems (cf. Example 1.5), for which there also exist specific algorithms that solve them efficiently [59].
@@
Continuity:
Under the assumption that MFCQ holds at $(x,u^{*}(x))$, the parametric solution $u^{*}$ is continuous [55, Theorem 5.3].
@@
Under the assumption that both MFCQ and CR hold at $(x,u^{*}(x))$, the parametric solution $u^{*}$ is locally Lipschitz [54, Theorem 3.6]. The same conclusion can be obtained if LICQ holds [53, Theorem 4.1] at $(x,u^{*}(x))$. We note also that since the satisfaction of LICQ implies the satisfaction of MFCQ and CR (cf. [54, Proposition 3.1]), [54, Theorem 3.6] is stronger than [53, Theorem 4.1].
@@
Under the assumptions of LICQ and SCS, the parametric solution $u^{*}$ is continuously differentiable [52, Theorem 2.1]. This last point was already noted in the optimization-based control literature in [41, Theorem 1]. In fact, if $f$ and $g$ belong to $\mathcal{C}^{p}(^{n},^{m})$, with $p\in\mathbb{Z}_{>0},p\geq 2$, the proof of [52, Theorem 2.1] can be adapted using the Implicit Function Theorem for higher degree of differentiability [60, Proposition 1B.5], to show that the parametric optimizer belongs to $\mathcal{C}^{p-1}(^{n})$.
@@
Similarly, if $f$ and $g$ are analytic in n, then the proof of [52, Theorem 2.1] can be adapted using the Analytic Function Theorem [61, Theorem 3.3.2], to show that the parametric optimizer is analytic in n.
If the constraint qualifications given above for the case of local Lipschitzness do not hold, the parametric optimizer can fail to be locally Lipschitz. To illustrate this, we revisit next an example due to Robinson [45].
@@
Problem (22) is a quadratic program with strongly convex objective function, smooth objective function and constraints, and for which Slater’s condition holds for every value of the parameter (this can be shown by noting that $\hat{u}=(0,0,2+|x_{2}|,0)$ satisfies all constraints strictly). Despite these nice properties, the parametric solution of (22) is not locally Lipschitz at $(x_{1},x_{2})=(0,0)$. Indeed, let $u^{*}:^{2}\to^{4}$ denote the parametric solution of (22), and $u_{4}^{*}:^{2}\to$ its fourth component, which is given by
@@
$u_{4}^{*}$ in Figure 1. For $s\geq 0$, let $p(s)=(s,s^{2}/2)$ and $q(s)=(s,0)$. Then, observe that $p(s)$ and $q(s)$ approach the origin as $s\to 0^{+}$, however,
$\frac{\left\lVert u_{4}^{*}(p(s))-u_{4}^{*}(q(s))\right\rVert}{\left\lVert p(s)-q(s)\right\rVert}=\frac{1}{s}.$
Since the right hand side of the previous expression can be made arbitrarily large by choose $s$ sufficiently small, it follows that $u^{*}$ is not locally Lipschitz at the origin. We also note that [54, Example 3.11] gives a similar example for a parametric quadratic program with a two-dimensional optimization variable, three-dimensional parameter, strongly convex objective function, smooth objective function and constraints, and for which Slater’s condition holds for every value of the parameter and the corresponding parametric optimizer also fails to be locally Lipschitz. $\bullet$
@@
(Regularity Properties of Parametric Optimizer): Suppose that $f$ and $g$ belong to $\mathcal{C}^{2}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $g(x,\cdot)$ is convex. Suppose that SC holds at $x_{0}\in^{n}$. Then,
@@
    if $f(x,\cdot)$ is strongly convex for all $x\in^{n}$, there exists a neighborhood $\mathcal{V}_{x_{0}}$ of $x_{0}$ such that $u^{*}$ is point-Lipschitz at $y$ for all $y\in\tilde{\mathcal{V}}_{x_{0}}$;
@@
    if $f(x,\cdot)$ is strongly convex for all $x\in^{n}$, $u^{*}$ has the Hölder property at $x_{0}$;
@@
    if $f(x,\cdot)$ is strictly convex for all $x\in^{n}$ and (1) has at least one minimizer for all $x\in^{n}$, $u^{*}$ is directionally differentiable at $x_{0}$.
@@
First we note that since in (i) and (ii), $f(x_{0},\cdot)$ is strongly convex, and in (iii) $f(x_{0},\cdot)$ is strictly convex and (1) has at least one minimizer at $x_{0}$, it follows that in all of (i)(ii)(iii), $u^{*}(x_{0})$ is unique and well-defined for all $x_{0}\in^{n}$.
To prove (i) we use [7, Theorem 6.4]. Since SC holds at $x_{0}$, by [62, Prop. 5.39], since $g(x_{0},\cdot)$ is convex, MFCQ holds at $(x_{0},u^{*}(x_{0}))$. Furthermore, since $f(x_{0},\cdot)$ is strongly convex and $g(x_{0},\cdot)$ is convex, the second-order condition SOC2 [7, Definition 6.1] holds (note that SOC2 is not guaranteed to hold if $f(x_{0},\cdot)$ is only strictly convex). All of this, together with the twice continuous differentiability of $f$ and $g$ imply, by [7, Theorem 6.4], that $u^{*}$ is point-Lipschitz at $x_{0}$. Now, since $g$ is continuous, there exists a neighborhood $\mathcal{V}_{x_{0}}$ of $x_{0}$ such that SC holds for all $y\in\mathcal{V}_{x_{0}}$. By repeating the same argument, $u^{*}$ is point-Lipschitz at $y$ for all $y\in\mathcal{V}_{x_{0}}$.
Now let us prove (ii). We use [63, Theorem 2.1], which gives a sufficient condition for the solution of a variational inequality to have the Hölder property. Recall that given a map $F:^{m}\to^{m}$, and a constraint set $\bar{\mathcal{C}}\subset^{m}$, a variational inequality refers to the problem of finding $u^{*}\in\bar{\mathcal{C}}$ such that $(u-u^{*})^{T}F(u^{*})\geq 0$ for all $u\in\bar{\mathcal{C}}$. For every fixed $x\in^{n}$, by taking the map $F$ to be the gradient of $f$ with respect to $u$ at $x$, and by taking $\bar{\mathcal{C}}$ to be the constraint set of (1) at $x$, a constrained optimization problem of the form (1) can be posed as a variational inequality, cf. [64]. Since $f$ is twice continuously differentiable and $f(x_{0},\cdot)$ is strongly convex, conditions (2.1) and (2.2) in [63, Theorem 2.1] hold. Note that condition (2.2) in [63, Theorem 2.1] is not guaranteed to hold if $f(x_{0},\cdot)$ is only strictly convex. Moreover, since MFCQ holds at $(x_{0},u^{*}(x_{0}))$ (because SC holds), by [65, Remark 3.6] the constraint set is pseudo-Lipschitzian [63, Definition 1.1]. All of this implies by [63, Theorem 2.1] that $u^{*}$ has the Hölder property at $x_{0}$.
Finally, (iii) follows from the fact that SC implies MFCQ and [66, Theorem 1]. Note that in this case, the assumptions of [66, Theorem 1] are satisfied by only requiring that $f(x,\cdot)$ is strictly convex for all $x\in^{n}$, instead of strongly convex for all $x\in^{n}$. ∎
@@
Figure 1: Surface plot of $u_{4}$, which is the fourth component of the parametric optimizer of Robinson’s counterexample, cf. (22). The purple solid and orange dashed lines correspond to the plot of $u_{4}(p(s))$ and $u_{4}(q(s))$ (defined in Example 3.1) respectively, for $s\geq 0$. The solid black lines connect points corresponding to $p(s)$ and $q(s)$. The plot shows that $u_{4}$ is continuous at the origin, in agreement with Proposition 3.2. However, since the slope of the line connecting $u_{4}(p(s))$ and $u_{4}(q(s))$ becomes arbitrarily large as $p$ and $q$ approach the origin, $u_{4}$ is not locally Lipschitz at the origin.
@@
(Scalar parametric quadratic programs have locally Lipschitz optimizers): Suppose that $f\in\mathcal{C}^{2}(^{n}\times^{1})$ and and for $i\in\{1,\ldots,p\}$, let $g_{i}^{0}:^{n}\to$, $g_{i}^{1}:^{n}\to$ belong to $\mathcal{C}^{2}(^{n})$ and